OpenStudy (anonymous):

How will I know if the intersection of two planes is a line or a plane if there's an intersection between them??

OpenStudy (precal):

if you are given a systems of equations, you will have a solution

OpenStudy (precal):

if not, then you would be given a picture to determine

OpenStudy (precal):

kind of hard to draw in 3d

OpenStudy (anonymous):

like for example this: \[x+y+z=1\]\[2x+2y+2z=2\] I know that they will intersect but how will i know if their intersection is plane or a line:?

OpenStudy (anonymous):

Two planes will intersect as a plane only if they are lying right next to each other... i.e. they are the same plane. So if you have one equation that is a multiple of another, like you do here, you know their intersection is a plane: the plane x+y+z=1.

OpenStudy (anonymous):

how about if the planes will have a line as their intersection? can you please give an example??thanks!

OpenStudy (anonymous):

Something like: x+y+z=1, 2x+y+2z=2, where one equation is not a multiple of the other.

OpenStudy (anonymous):

http://www.youtube.com/watch?v=jozabh0lFmo

OpenStudy (anonymous):

aahhh.. ok!, thanks a lot:)

OpenStudy (anonymous):

And there is of course the final case: e.g. x+y+z=1, x+y+z=2, where you get an inconsistency and the planes do not intersect at all (as they are parallel). You're welcome! :)

OpenStudy (anonymous):

@cinar i already checked that out, it helped me a little bit, thanks by the way. @Erin001001 yes ☺